# Unsupervised Meta-Learning via In-Context Learning

### v2

@@ 4.1 Datasets and baselines
For the evaluation, we use two generic object recognition datasets, i.e., miniImageNet [39] and CIFAR-fs [2], and three fine-grained image classification datasets, i.e., CUB [52], Aircraft [31], and Meta-iNat [55]. While miniImageNet and CIFAR-fs share some similarities with ImageNet-1k, as some classes appears in all three datasets, CUB, Aircraft, and Meta-iNat focus on more specialized domains, ensuring a rigourous cross-domain evaluation. Each dataset is split into training, validation, and test sets following the splits in [39] and [2] for miniImageNet and CIFAR-fs, respectively, and in [45] and [34] for the remaining datasets. All labels are removed from the datasets during the training phase.
We compare the results of CAMeLU with standard UML approaches such as CACTUs [19], UMTRA [21], Meta-GMVAE [26], and PsCo [20]. These methods are evaluated in-domain as recommended in the original papers, with training and testing performed on the same dataset. While this setup is relatively simpler than the cross-domain evaluation employed for CAMeLU, applying these methods in a cross-domain scenario may not be fair, as they were not explicitly designed for such a challenging scenario. Only PsCo [20] is further evaluated in a cross-domain setting, as the authors demonstrate its adaptability to this scenario through an additional adaptation phase to the test domain. We also compare CAMeLU with BECLR [34] in Section 4.5, a contrastive framework for unsupervised few-shot learning. Although BECLR does not utilize meta-learning, it serves as a strong baseline for unsupervised few-shot learning. Additionally, CAMeLU is compared to CAML [12], a supervised meta-learning approach that assumes tasks are available both during the training and testing phases and leverages the in-context ability of transformer architectures to generalize to new tasks. Due to their similarities, we refer to CAML as the supervised counterpart of CAMeLU. Furthermore, Section 4.6 provides a comparative analysis with two fine-tuned state-of-the-art self-supervised trained networks, namely SimCLR [7] and SwAV [6].
@@ 4.2 Training details
We report the results following the $N$-way $K$-shot classification task typical of meta-learning algorithms, where $N=5$ and $K=1$ or $K=5$. All models are trained for $100$ epochs, $500$ episodes per epoch, and $100$ fine-tuning steps at test time if required by the method. For CAMeLU, we choose not to apply any fine-tuning strategy to prove the strength of the training stage that does not require additional parameter updates during the inference phase. Furthermore, we introduce ImageNet-964, a variant of the original ImageNet-1k dataset [8] where the classes belonging to the validation and test splits of miniImageNet are removed to prevent data leakage when evaluating the model on miniImageNet - a problem that is not taken into consideration by previous studies [12, 20]. To provide a fair comparison, all cross-domain methods are trained using ImageNet-964. For CAMeLU, we use a feature extractor pre-trained using ResNet-50 [18] on ImageNet-964 and a class encoder consisting of a single learnable layer. In Appendix A.4, we also report the results with different feature extractors. The transformer encoder consists of 8 encoder layers, each incorporating a multi-head self-attention block with 8 attention heads, an MLP with a reverse bottleneck of $3072$ and an input-output feature size of $2304$, where the last $2048$ features is the ResNet-50 feature vector appended to the $256$ encoding of the labels. Finally, a single projection layer maps the transformer output to the predicted category. The model is trained with the Adam optimizer with a learning rate of $10^{-5}$ and a warmup cosine scheduler as proposed in the original paper [48]. To account for statistical variations, each algorithm is run three times in full and the complete results reporting the standard deviations are presented in Appendix A.10. The experiments are executed using Python and the PyTorch library on an Nvidia GeForce RTX 3070 Ti Laptop GPU with 8GB of VRAM. Further ablation experiments and competitors required a more capable architecture and are executed on an Nvidia A100-SXM4 GPU with 40GB of VRAM. More details about the training settings can be found in Appendix A.1, while the code will be released upon acceptance of the paper.
@@ 4.3 Comparative results
Table 1: Performance comparison on miniImageNet, CIFAR-fs, CUB, Aircraft, and Meta-iNat datasets for 5-way 1-shot and 5-way 5-shot scenarios. Cross-domain approaches are trained using ImageNet-964 and a ResNet-50 feature extractor. The symbol $\dagger$ indicates results that are affected by data leakage. The bold font highlights the best performing UML approach for each setting. Results show the average across three complete runs of the algorithms. Complete results with standard deviations are reported in Table 10 in Appendix A.10.
@@
Table 1 provides an overview of the experimental results for both the 5-way 1-shot and the 5-way 5-shot scenarios. The results demonstrate that CAMeLU outperforms the existing UML methods, regardless of the difference in the evaluation setting. As highlighted in Section 4.1, CACTUs, UMTRA, and Meta-GMVAE are evaluated only in-domain, requiring knowledge about the test domain prior to training. This is not necessary for CAMeLU as it demonstrates high performance in the challenging cross-domain scenario. Even compared to PsCo, the only UML method designed for cross-domain applications, CAMeLU exhibits a performance improvement in all datasets. Furthermore, PsCo requires a fine-tuning phase to adapt to the test domain, whereas CAMeLU achieves good performance with a single forward pass, enhancing its applicability to real-time applications. It is also worth noting that CAMeLU achieves comparable performance to its supervised counterpart, CAML, when evaluated on miniImageNet and it even outperforms CAML when evaluated on more dissimilar domains, such as CUB, Aircraft, and Meta-iNat. This finding highlights the efficacy of the task construction strategy used in CAMeLU, which acts as a sort of task augmentation and enhances the generalization capability of the model.
@@ 4.4 Memorization to generalization phase shift
During the training of CAMeLU, we observed a distinct trend in the validation accuracy, similar to the findings in [23]. Figure 2 illustrates this pattern, showing the validation accuracy relative to its initial value, or, in other words, how much the model learns from datasets different from the one we are training on. Specifically, the curves in Figure 2 resemble a logistic curve,
Figure 2: Analysis of learning behavior when transferring knowledge from different prior datasets. The relative validation accuracy shows the difference between the current and first epoch accuracy on the validation set of miniImageNet, CIFAR-fs, CUB, and Aircraft. CAMeLU is trained with ImageNet-964.
which can be divided into three phases that we denote as memorization, learning, and generalization. In the memorization phase, the model memorizes the tasks seen during training and extends this knowledge to unseen tasks, resulting in a slight improvement for datasets with high similarity with ImageNet-964 (e.g., miniImageNet and CIFAR-fs). For the other datasets, instead, transferring this knowledge can even result in a performance decrease due to the intrinsic domain distance of the dataset (see CUB and Aircraft, which are fine-grained datasets). As training progresses and the model observes more tasks, the learning phase occurs. This phase is characterized by a transition to the learning-to-learn state where the model learns to identify the tasks and to extract the features that are more useful for solving them. The duration of this phase varies, with datasets like miniImageNet and CIFAR-fs exhibiting rapid learning within approximately $10$ epochs, while datasets such as CUB and Aircraft may necessitate up to 40 epochs. This timespan depends on several factors, including the similarity between the training and evaluation datasets, the size of the test dataset, and the model’s learning ability [23, 35]. For instance, CUB, with its fine-grained nature and small test set size (around $1770$ images), necessitates a longer learning phase compared to the miniImageNet dataset (which has a test set with 12 0001200012\,00012 000 images). Subsequently, in the generalization phase, the model can generalize to tasks significantly different from those observed during training using a single forward pass. Further analyses about the generalization capabilities of CAMeLU and the number of epochs required for reaching the generalization phase are presented in Section 4.5 and Appendix A.8.
@@ 4.5 Generalization on small-scale datasets
Table 2: Accuracy results obtained training PsCo, BECLR, and CAMeLU with a small-scale dataset, namely miniImageNet, denoted as (mini) in the table. Results show both in-domain performance (on the test set of miniImageNet) and cross-domain performance on CIFAR-fs, CUB, Aircraft, and Meta-iNat. The average results across three complete runs of the algorithms are reported. Complete results with standard deviations are presented in Table 11 in Appendix A.10.
@@
While most studies on training transformer architectures focus on large-scale training datasets, we investigate the generalization capabilities of CAMeLU using a small-scale training dataset. Specifically, we train CAMeLU on miniImageNet and evaluate its performance both in-domain (i.e., on the test set of miniImageNet) and cross-domain on CIFAR-fs, CUB, Aircraft, and Meta-iNat. CAMeLU demonstrates effective generalization in this scenario, showing impressive performance in both in-domain and cross-domain settings, as shown in Table 2, surpassing PsCo and BECLR by a large margin.
Additionally, a comparison of CAMeLU’s performance when trained on a small-scale dataset like miniImageNet (Table 2), on ImageNet-964 (Table 1), and on a large-scale dataset (Table 5 in Appendix A.3) show that our method is only slightly affected by the size of the training dataset.
Figure 3: Relative validation accuracy comparison between CAMeLU (orange) and CAML (blue) when evaluated in-domain on miniImagenet. The relative validation accuracy is computed as described in Section 4.4. The curve obtained with CAMeLU reflects the three phases of memorization, learning, and generalization even when using a small-scale dataset.
This robustness enhances CAMeLU’s applicability to scenarios where only a small unlabeled training dataset is available, which is common in real-world applications.
Finally, Figure 3 shows the relative validation accuracy of CAMeLU and CAML while trained and evaluated on miniImageNet. While the curve obtained with CAMeLU reflects the three phases of memorization, learning, and generalization discussed in Section 4.4, the relative validation accuracy of CAML remains flat. This difference may be attributed to the task creation mechanism of CAMeLU, which acts as a task augmentation strategy, increasing the variability of tasks presented to the model during training and thereby enhancing its generalization capabilities.
@@ 4.6 Comparison with SSL methods
Table 3: Comparison between CAMeLU and SSL approaches for the 5-way 1-shot and 5-way 5-shot scenario on miniImageNet, CIFAR-fs, CUB, Aircraft, and Meta-iNat. The symbol $\dagger$ indicates results that are affected by data leakage. Results show the average across three complete runs of the algorithms. Complete results with standard deviations are reported in Table 12 in Appendix A.10.
@@
In this section, we compare CAMeLU with SimCLR [7] and SwAV [6]. For training SSL methods in our experiments, we employed a backbone network with a ResNet-50 architecture pre-trained on ImageNet-1k obtained from PyTorch Lightning Bolts [3]. While this setup leads to data leakage when evaluated on miniImageNet, due to overlap between the test and training classes, pre-training these SSL approaches from scratch using a different training dataset was beyond our available computational resources. To facilitate model adaptation to the test domain, we fine-tuned a classification layer on top of the pre-trained backbone using SGD with initial learning rate of $0.01$, momentum of $0.9$, weight decay of $10^{-4}$, and $100$ fine-tuning steps per task, following [20]. This setup differs from the evaluation setting used in CAMeLU, where predictions are obtained with a single forward pass, leveraging the in-context learning ability of transformer architectures. However, SSL approaches must adapt to the test domain before label predictions, resulting in a less challenging evaluation setting than CAMeLU. Results for SSL approaches are averaged over $500$ test tasks and presented in Table 3. While SSL approaches outperform CAMeLU on miniImageNet and CIFAR-fs, their performance decreases when evaluated on the other datasets. CUB, Aircraft, and Meta-iNat are fine-grained datasets significantly different from ImageNet-1k, challenging the transferability of features learned by SSL methods to these datasets. Moreover, the high performance on miniImageNet and CIFAR-fs may be attributed to the presence of data leakage with ImageNet-1k and the high similarity with CIFAR-fs, as discussed in Section 4.2. CAMeLU, in contrast, demonstrates effective generalization to tasks sampled from these datasets, once again highlighting its generalization ability over mere memorization.
@@ A.1 Experimental details
For training CAMeLU, we use ImageNet-964, which is a variant of the original ImageNet-1k dataset [8] where classes belonging to the validation and test splits of miniImageNet [39] are removed. This results in a total of 1 234 48712344871\,234\,4871 234 487 images for training the model compared to the 1 281 16712811671\,281\,1671 281 167 in the original ImageNet-1k dataset. When a multi-dataset approach is utilized for training CAMeLU (see Appendix A.3), MSCOCO [28] and Fungi [40] are loaded into the program and used together with ImageNet-964 for creating the whole training dataset. MSCOCO is a dataset originally proposed for object detection, where each image is assigned to $G$ classes corresponding to the $G$ objects present in it. To use it for image classification, we replicate each image $G$ times and we assign to each of them one of the $G$ classes. In this way, we obtain a dataset with 117 266117266117\,266117 266 images for training, and we rely on the fact that the transformer is capable of applying self-attention to the object of the class in question. Fungi is a fine-grained dataset with a size of only 64 3076430764\,30764 307, which is two orders of magnitude smaller than the ImageNet-964 and MSCOCO. For evaluation and for in-domain training of the baselines, we use miniImageNet, CIFAR-fs, CUB, Aircraft, and Meta-iNat. miniImageNet is split into train/validation/test using the splits proposed in [39], resulting in 38 4003840038\,40038 400 images for training, $9600$ for validation, and 12 0001200012\,00012 000 for testing. The same number of images are also present in CIFAR-fs, and the splits follow the work in [2]. CUB and Aircraft, instead, are two fine-grained datasets with a smaller size compared to the others. CUB [52] consists of $8239$ in the training set, $1779$ in the validation set, and $1770$ in the test set, while Aircraft has respectively $7000$/$1500$/$1500$ images in the train/validation/test sets [45]. Finally, Meta-iNat [55] consists of 243 986243986243\,986243 986 images split into $1135$ classes, with $227$ reserved for testing. All images are resized to $224\times 224$ and normalized with zero mean and unit variance before input into the model. For the UML baselines, due to the smaller model size utilized in the esperiments, images are resized to $84\times 84$, as suggested in the original papers [21, 19, 26, 20].
@@
The architecture used for CAMeLU consists of a fixed pre-trained feature extractor, a class encoder, and a transformer encoder. The feature extractor is pre-trained using ResNet-50 on ImageNet-964 following the same architecture and hyperparameters in [18]. The class encoder is a single learnable layer with a dimensionality of $256$ and initialized with Kaiming initialization [17]. Image embeddings are concatenated with class embeddings before being fed into the transformer encoder. This results in a vector with a total length of $2304$, composed of $2048$ features from the image embeddings and $256$ from the label embeddings. When ablating the feature extractor with CLIP [36] in Appendix A.4, a ViT-B/16 encoder architecture is utilized and downloaded from the Hugging Face website [57]. The pre-training is performed using a large dataset with $400$ million (image, text) pairs [36], and it is fixed during the training phase of CAMeLU. The image embedding’s output size decreases to $1024$, resulting in an input vector length of $768$ and a reduced memory complexity compared to ResNet-50. The transformer encoder comprises 8 encoder layers. Each layer consists of 8 attention heads and an MLP with a reversed bottleneck of $3072$ (with GeLU activation function). A projection layer completes the model architecture to map the transformer output to a class prediction. This architecture enables us to store the entire model in an Nvidia GeForce RTX 3070 Ti Laptop GPU with 8GB of VRAM, while a further reduction of memory can be achieved by utilizing CLIP on ViT-B/16 as feature extractor, which requires only 4 GB of VRAM to store the entire model.
For the task creation mechanism, 3 augmentation functions are selected from a list comprising cropping, rotation, horizontal flip, grayscale, color jittering, gaussian blur, and random affine transformation. The exact parameters used in our experiments for each augmentation function are detailed in Table 4. For the query set, an additional pixel-level mixing strategy with $\lambda\sim Beta(\alpha,\beta)$ with $\alpha=1,\beta=1$ and $\lambda\in(0,0.5)$ is utilized. More details about this selection choice can be found in Appendix A.6.

### v3

@@ 4.1 Datasets and baselines
For the evaluation, we use two generic object recognition datasets, i.e., miniImageNet (Ravi & Larochelle, 2016) and CIFAR-fs (Bertinetto et al., 2019), and three fine-grained image classification datasets, i.e., CUB (Wah et al., 2011), Aircraft (Maji et al., 2013), and Meta-iNat (Wertheimer & Hariharan, 2019). While miniImageNet and CIFAR-fs share some classes with ImageNet-1k, CUB, Aircraft, and Meta-iNat focus on more specialized domains, ensuring a rigorous cross-domain evaluation. Each dataset is split into training, validation, and test sets following the splits in Ravi & Larochelle (2016) and Bertinetto et al. (2019) for miniImageNet and CIFAR-fs, respectively, and in Triantafillou et al. (2019) and Poulakakis-Daktylidis & Jamali-Rad (2024) for the remaining datasets. All labels are removed from the datasets during the training phase.
We compare CAMeLU with standard UML approaches such as CACTUs (Hsu et al., 2018), UMTRA (Khodadadeh et al., 2019), Meta-GMVAE (Lee et al., 2020), and PsCo (Jang et al., 2022). These methods are evaluated in-domain as recommended in the original papers, with training and testing performed on the same dataset. While this setup is relatively simpler than the cross-domain evaluation employed for CAMeLU, applying these methods in a cross-domain scenario may not be fair, as they were not explicitly designed for such a challenging scenario. Only PsCo (Jang et al., 2022) is further evaluated in a cross-domain setting, as the authors demonstrate its adaptability to this scenario through an additional adaptation phase to the test domain. We also compare CAMeLU with BECLR (Poulakakis-Daktylidis & Jamali-Rad, 2024) in Sect. 4.5, a contrastive framework for unsupervised few-shot learning, and CAML (Fifty et al., 2024), a supervised meta-learning approach that assumes tasks are available both during the training and testing phases and leverages the in-context ability of transformer architectures to generalize to new tasks. Due to their similarities, we refer to CAML as the supervised counterpart of CAMeLU. Furthermore, Sect. 4.6 provides a comparative analysis with two fine-tuned state-of-the-art self-supervised trained networks, namely SimCLR (Chen et al., 2020) and SwAV (Caron et al., 2020).
@@ 4.2 Training details
We report the results following the $N$-way $K$-shot classification task typical of meta-learning algorithms, where $N=5$ and $K=1$ or $K=5$. All models are trained for $100$ epochs with $500$ episodes per epoch. Fine-tuning at test time (100 steps) is applied only if required. For CAMeLU, we do not apply any fine-tuning step to demonstrate the strength of its training stage, which does not require additional parameter updates during inference. Furthermore, we introduce ImageNet-964, a variant of ImageNet-1k (Deng et al., 2009) where classes from the validation and test splits of miniImageNet are removed to prevent data leakage—a problem that is not taken into consideration by previous studies (Fifty et al., 2024; Jang et al., 2022). To provide a fair comparison, all cross-domain methods are trained on ImageNet-964. For CAMeLU, we use a ResNet-50 (He et al., 2016) feature extractor pre-trained on ImageNet-964 and a class encoder that maps one-hot label vectors to a $256$-dimensional space. In Appendix A.4, we also report the results with different feature extractors. The transformer encoder consists of 8 layers, each with an eight-head self-attention block, an MLP, and a single projection layer that maps the transformer output to the predicted category. The model is trained with the Adam optimizer with a learning rate of $10^{-5}$ and a warmup cosine scheduler (Vaswani et al., 2017). To account for statistical variations, each algorithm is run three times in full, and the complete results reporting the standard deviations are presented in Appendix A.10. The experiments are executed using Python and the PyTorch library on an Nvidia GeForce RTX 3070 Ti Laptop GPU with 8GB of VRAM, while ablation studies and competitors are executed on an Nvidia A100-SXM4 GPU with 40GB of VRAM. More details about the training settings can be found in Appendix A.1, and the code is available at https://github.com/bracca95/CAMeLU.git.
@@ 4.3 Comparative results
Table 1: Performance comparison on miniImageNet, CIFAR-fs, CUB, Aircraft, and Meta-iNat datasets for 5-way 1-shot and 5-way 5-shot scenarios. Cross-domain approaches are trained using ImageNet-964 and a ResNet-50 feature extractor. The symbol $\dagger$ indicates results that are affected by data leakage. The bold font highlights the best performing UML approach for each setting. Results show the average across three complete runs of the algorithms. Complete results with standard deviations are reported in Tab. 12 in Appendix A.10.
@@
Table 1 provides an overview of the experimental results for both the 5-way 1-shot and the 5-way 5-shot scenarios. The results demonstrate that CAMeLU outperforms the existing UML methods, regardless of the difference in the evaluation setting. As highlighted in Sect. 4.1, CACTUs, UMTRA, and Meta-GMVAE are evaluated only in-domain, requiring knowledge about the test domain prior to training. This is not necessary for CAMeLU as it demonstrates high performance in the challenging cross-domain scenario. Even compared to PsCo, the only UML method designed for cross-domain applications, CAMeLU exhibits a performance improvement across all datasets. Furthermore, PsCo requires a fine-tuning phase to adapt to the test domain, whereas CAMeLU achieves good performance with a single forward pass, enhancing its applicability to real-time applications. It is also worth noting that CAMeLU achieves comparable performance to its supervised counterpart, CAML, when evaluated on miniImageNet and it even outperforms CAML when evaluated on more dissimilar domains, such as CUB, Aircraft, and Meta-iNat. This finding highlights the efficacy of the task construction strategy used in CAMeLU, which acts as a sort of task augmentation and enhances the generalization capability of the model.
@@ 4.4 Memorization to generalization phase shift
Figure 3: Analysis of learning behavior when transferring knowledge from a different prior dataset. The relative validation accuracy shows the difference between the current and first epoch accuracy on the validation set of miniImageNet, CIFAR-fs, CUB, and Aircraft. CAMeLU is trained with ImageNet-964.
During the training of CAMeLU, we observed a distinct trend in the validation accuracy, similar to the findings in Kirsch et al. (2022). Fig. 3 illustrates this pattern, showing the validation accuracy relative to its initial value, or, in other words, how much the model learns from datasets different from the one we are training on. Specifically, the curves in Fig. 3 resemble a logistic curve, which can be divided into three phases that we denote as memorization, learning, and generalization. In the memorization phase, the model memorizes the tasks seen during training and extends this knowledge to unseen tasks, resulting in a slight improvement for datasets with high similarity with ImageNet-964 (e.g., miniImageNet and CIFAR-fs). For the other datasets, instead, transferring this knowledge can even result in a performance decrease due to the intrinsic domain distance of the dataset (see CUB and Aircraft, which are fine-grained datasets). As training progresses and the model observes more tasks, the learning phase occurs. This phase is characterized by a transition to the learning-to-learn state where the model learns to identify the tasks and to extract the features that are more useful for solving them. The duration of this phase varies, with datasets like miniImageNet and CIFAR-fs exhibiting rapid learning within approximately $10$ epochs, while datasets such as CUB and Aircraft may necessitate up to 40 epochs. This timespan depends on several factors, including the similarity between the training and evaluation datasets, the size of the test dataset, and the model’s learning ability (Kirsch et al., 2022; Power et al., 2022). For instance, CUB, with its fine-grained nature and small test set size (around $1770$ images), necessitates a longer learning phase compared to the miniImageNet dataset (which has a test set with 12 0001200012\,00012 000 images). Subsequently, in the generalization phase, the model can generalize to tasks significantly different from those observed during training using a single forward pass. Further analyses about the generalization capabilities of CAMeLU and the number of epochs required for reaching the generalization phase are presented in Sect. 4.5 and Appendix A.8.
@@ 4.5 Generalization on small-scale datasets
Table 2: Accuracy results obtained training PsCo, BECLR, and CAMeLU with a small-scale dataset, namely miniImageNet, denoted as (mini) in the table. Results show both in-domain performance (on the test set of miniImageNet) and cross-domain performance on CIFAR-fs, CUB, Aircraft, and Meta-iNat. The average results across three complete runs of the algorithms are reported. Complete results with standard deviations are presented in Tab. 13 in Appendix A.10.
@@
While most studies on training transformer architectures focus on large-scale training datasets, we investigate the generalization capabilities of CAMeLU using a small-scale training dataset. Specifically, we train CAMeLU on miniImageNet and evaluate its performance both in-domain (i.e., on the test set of miniImageNet) and cross-domain on CIFAR-fs, CUB, Aircraft, and Meta-iNat. CAMeLU demonstrates effective generalization in this scenario, showing impressive performance in both in-domain and cross-domain settings, as shown in Tab. 2, surpassing PsCo and BECLR by a large margin.
Additionally, a comparison of CAMeLU’s performance when trained on a small-scale dataset like miniImageNet (Tab. 2), on ImageNet-964 (Tab. 1), and on a large-scale dataset (Tab. 5 in Appendix
Figure 4: Relative validation accuracy of CAMeLU (orange) and CAML (blue) when evaluated in-domain on miniImagenet and computed as in Sect. 4.4. The curve obtained with CAMeLU reflects the three phases of memorization, learning, and generalization even when using a small-scale dataset.
A.3) show that our method is only slightly affected by the size of the training dataset. This robustness enhances CAMeLU’s applicability to scenarios where only a small unlabeled training dataset is available, which is common in real-world applications.
Finally, Fig. 4 shows the relative validation accuracy of CAMeLU and CAML while trained and evaluated on miniImageNet. While the curve obtained with CAMeLU reflects the three phases of memorization, learning, and generalization discussed in Sect. 4.4, the relative validation accuracy of CAML remains flat. This difference may be attributed to the task creation mechanism of CAMeLU, which acts as a task augmentation strategy, increasing the variability of tasks presented to the model during training and thereby enhancing its generalization capabilities.
@@ 4.6 Comparison with SSL methods
Table 3: Comparison between CAMeLU and SSL approaches for the 5-way 1-shot and 5-way 5-shot scenario on miniImageNet, CIFAR-fs, CUB, Aircraft, and Meta-iNat. The symbol $\dagger$ indicates results that are affected by data leakage. Results show the average across three complete runs of the algorithms. Complete results with standard deviations are reported in Tab. 14 in Appendix A.10.
@@
In this section, we compare CAMeLU with SimCLR (Chen et al., 2020) and SwAV (Caron et al., 2020). For training SSL methods in our experiments, we employed a backbone network with a ResNet-50 architecture pre-trained on ImageNet-1k and obtained from PyTorch Lightning Bolts (Borovec et al., 2022). While this setup leads to data leakage when evaluated on miniImageNet, due to overlap between the test and training classes, pre-training these SSL approaches from scratch using a different training dataset was beyond our available computational resources. To facilitate model adaptation to the test domain, we fine-tuned a classification layer on top of the pre-trained backbone using SGD with an initial learning rate of $0.01$, momentum of $0.9$, weight decay of $10^{-4}$, and $100$ fine-tuning steps per task, following Jang et al. (2022). This setup differs from the evaluation setting used in CAMeLU, where predictions are obtained with a single forward pass, leveraging the in-context learning ability of transformer architectures. However, SSL approaches must adapt to the test domain before label predictions, resulting in a less challenging evaluation setting than CAMeLU. Results for SSL approaches are averaged over $500$ test tasks and presented in Tab. 3. While SSL approaches outperform CAMeLU on miniImageNet and CIFAR-fs, their performance decreases when evaluated on the other datasets. CUB, Aircraft, and Meta-iNat are fine-grained datasets significantly different from ImageNet-1k, challenging the transferability of features learned by SSL methods to these datasets. Moreover, the high performance on miniImageNet and CIFAR-fs may be attributed to the presence of data leakage with ImageNet-1k and the high similarity with CIFAR-fs, as discussed in Sect. 4.2. CAMeLU, in contrast, demonstrates effective generalization to tasks sampled from these datasets, once again highlighting its generalization ability over mere memorization.
@@ 5 Conclusion
Ethics statement and reproducibility guidelines
In this work, we used well-established, publicly available datasets to train and evaluate our architecture. While these datasets and pre-trained models provide a valuable foundation for research, we acknowledge the potential for inherent biases that may not fully represent diverse real-world scenarios. We have taken every precaution to ensure that our experiments are conducted responsibly, with no intention of causing harm or perpetuating any biases present in the data. Furthermore, we declare no conflicts of interest in the execution or reporting of this research. Our objective is to present the findings in a transparent manner and contribute positively to the broader research community.
To ensure the reproducibility of our experiments, we have provided the code and detailed instructions on how to run the experiments. The general configuration of our model is described in Sect. 4, with additional technical details outlined in Sect. A.1 of the Appendix. Moreover, we have employed random seed initialization to ensure consistency across runs. The complete codebase, models, and pre-trained weights are available on GitHub  [^1] https://github.com/bracca95/CAMeLU.git to facilitate further research and replication.
Acknowledgments
This work was supported by the “Knowledge Foundation” (KK-stiftelsen).
@@ A.1 Experimental details
For training CAMeLU, we use ImageNet-964, which is a variant of the original ImageNet-1k dataset (Deng et al., 2009) where classes belonging to the validation and test splits of miniImageNet (Ravi & Larochelle, 2016) are removed. This results in a total of 1 234 48712344871\,234\,4871 234 487 images for training the model compared to the 1 281 16712811671\,281\,1671 281 167 in the original ImageNet-1k dataset. When a multi-dataset approach is utilized for training CAMeLU (see Appendix A.3), MSCOCO (Lin et al., 2014) and Fungi (Schroeder & Cui, 2018) are loaded into the program and used together with ImageNet-964 for creating the whole training dataset. MSCOCO is a dataset originally proposed for object detection, where each image is assigned to $G$ classes corresponding to the $G$ objects present in it. To use it for image classification, we replicate each image $G$ times and we assign to each of them one of the $G$ classes. In this way, we obtain a dataset with 117 266117266117\,266117 266 images for training, and we rely on the fact that the transformer is capable of applying self-attention to the object of the class in question. Fungi is a fine-grained dataset with a size of only 64 3076430764\,30764 307, which is two orders of magnitude smaller than ImageNet-964 and MSCOCO. For evaluation and for in-domain training of the baselines, we use miniImageNet, CIFAR-fs, CUB, Aircraft, and Meta-iNat. miniImageNet is split into train/validation/test using the splits proposed in Ravi & Larochelle (2016), resulting in 38 4003840038\,40038 400 images for training, $9600$ for validation, and 12 0001200012\,00012 000 for testing. The same number of images are also present in CIFAR-fs, and the splits follow the work in Bertinetto et al. (2019). CUB and Aircraft, instead, are two fine-grained datasets with a smaller size compared to the others. CUB (Wah et al., 2011) consists of $8239$ in the training set, $1779$ in the validation set, and $1770$ in the test set, while Aircraft has respectively $7000$/$1500$/$1500$ images in the train/validation/test sets (Triantafillou et al., 2019). Finally, Meta-iNat (Wertheimer & Hariharan, 2019) consists of 243 986243986243\,986243 986 images split into $1135$ classes, with $227$ reserved for testing. All images are resized to $224\times 224$ and normalized with zero mean and unit variance before input into the model. For the UML baselines, due to the smaller model size utilized in the experiments, images are resized to $84\times 84$, as suggested in the original papers (Khodadadeh et al., 2019; Hsu et al., 2018; Lee et al., 2020; Jang et al., 2022).
@@
The architecture used for CAMeLU consists of a fixed pre-trained feature extractor, a class encoder, and a transformer encoder. The feature extractor is pre-trained using ResNet-50 on ImageNet-964 following the same architecture and hyperparameters in He et al. (2016). The class encoder is a single learnable layer with a dimensionality of $256$ and initialized with Kaiming initialization (He et al., 2015). Image embeddings are concatenated with class embeddings before being fed into the transformer encoder. This results in a vector with a total length of $2304$, composed of $2048$ features from the image embeddings and $256$ from the label embeddings. When ablating the feature extractor with CLIP (Radford et al., 2021) in Appendix A.4, a ViT-B/16 encoder architecture is utilized and downloaded from the Hugging Face website (Wolf et al., 2019). The pre-training is performed using a large dataset with $400$ million (image, text) pairs (Radford et al., 2021), and it is fixed during the training phase of CAMeLU. The output size of the image embedding is reduced to $1024$, with $768$ features from the image embedding, which results in a reduced memory complexity compared to ResNet-50. The transformer encoder comprises 8 encoder layers. Each layer consists of 8 attention heads and an MLP with a reversed bottleneck of $3072$ (with GeLU activation function). A projection layer completes the model architecture to map the transformer output to a class prediction. This architecture enables us to store the entire model in an Nvidia GeForce RTX 3070 Ti Laptop GPU with 8GB of VRAM, while a further reduction of memory can be achieved by utilizing CLIP on ViT-B/16 as feature extractor, which requires only 4 GB of VRAM to store the entire model.
For the task creation mechanism, 3 augmentation functions are selected from a list comprising cropping, rotation, horizontal flip, grayscale, color jittering, gaussian blur, and random affine transformation. The exact parameters used in our experiments for each augmentation function are detailed in Tab. 4. For the query set, an additional pixel-level mixing strategy with $\lambda\sim Beta(\alpha,\beta)$ with $\alpha=1,\beta=1$ and $\lambda\in(0,0.5)$ is utilized. More details about this selection choice can be found in Appendix A.6.